Grade 11Physical Chemistry

the rate of diffusion of a gas is 2.92 times greater than that of ammonia gas . determine the molecular weight of the gas

Arvind Singh
7 Years agoGrade 11

1 Answer

Manda Snehashish Reddy
7 Years ago
Let molecular weight of unknown gas be M. We know that r1/r2 =√M2/√M1. Here M2= molecular weight of ammonia gas, which is 17amu. 2.92=√17/√M. Solving it M= 1.9938074 (or) M= 2